The ratio of their ages of Elton and David is 3:5. In 6 years time the ratio of their ages will be 3:4. How old is David?

Answer by dkppathak(439)

The ratio of their ages of Elton and David is 3:5. In 6 years time the ratio of their ages will be 3:4. How old is David?
let present ratio of age ofelton and david age is =3x/5x
as per given conditions after 6 yrs
3x+6/5x+=3/4
4(3x+6)=3(5x+6)
12x + 24=15x +18
12x-15x=18-24
-3x=-6
x=2
age of elton is 3x=3x2=6yrs
age of david is 5x=5x2=10


Answer David age present is 10 yrs